OpenStudy (anonymous):

A tree of height 50 feet casts a shadow 80 feet long at a certain time of day. A second tree near to the first casts a shadow 100 feet long at the same time. How many feet taller is the second tree than the first?

OpenStudy (anonymous):

50/80 = x/100 62.5 = x 62.5 feet tall